OpenStudy (czesc):

INSTANT FAN/MEDAL What degree of the Maclaurin polynomial is required so that the error in the approximation of e^0.6 is less than 0.001? (no idea how to start)

OpenStudy (shirley27):

What do you know about the different error tests that can be used with the series?

OpenStudy (czesc):

@Shirley27 Yes, which do you suggest I apply here and how?

OpenStudy (czesc):

thanks for responding by the way, I appreciate it !

OpenStudy (shirley27):

no problem :) I would suggest using the langrange error bound in this case since the maclaurin series isn't alternating

OpenStudy (shirley27):

the actual langrange error bound is through this formula\[error \le \frac{ M \times x ^{n+1} }{ (n+1)! }\]

OpenStudy (shirley27):

where \[M \ge f ^{(n+1)}(x)\] and x is a number between 0 and the estimated number that produces the largest derivative

OpenStudy (reemii):

As @Shirley27 mentioned, take the formula of the rest: \[\frac{f^{(n+1)}(\xi)x^{n+1}}{(n+1)!} \qquad (*)\] We do not know the value of \(\xi \in (0,x)\). Therefore we look for an upper bound of (*). Since \(f^{(n+1)}(x)\) is an increasing function (check it if necessary), an upper bound of (*) is found at point \(x\): \(f^{(n+1)}(x)x^{n+1}/(n+1)!\) is an upper bound. (\(x=0.6\)). (If \(f^{(n+1)}\) is a positive decreasing function, the upper bound is found by setting \(\xi=0\).)

OpenStudy (reemii):

An upper bound of (*) is \(\frac{e^{0.6} (0.6)^{n+1}}{(n+1)!}\), for all \(n\). You just have to evaluate this until you obtain a value smaller than 0.001.